# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 6.2. Defense Against Privacy Attacks on LLMs
Defense Against Gradient Leakage Attack. There are several mitigation strategies to defend against those gradient-based attack methods, e.g., random noise insertion (wei2020framework, ), differential privacy (geyer2017differentially, ), and homomorphic encryption (aono2017privacy, ). Building upon prior research on vision model attacks (zhu2019deep, ), (wei2020framework, ), the defense mechanisms involving the addition of Gaussian or Laplacian noise to gradients and DP-SGD coupled with additional clipping (abadi2016deep, ) can form an effective defense against gradient leakage attacks. However, it may sacrifice the model’s utility to a certain extent. Prior works explored various techniques, such as (wu2023learning, ), (raeini2023privacy, ), and (huang2022large, ) to defend against gradient leakage attacks in the language domain for small NLP models. As these existing defense mechanisms are not tailored for LLMs, further studies are needed to develop defense mechanisms against gradient leakage attacks on LLMs.
@@
Defense Against Membership Inference Attack. In order to mitigate MIA in the language domain, several mechanisms are proposed, including dropout, model stacking, differential privacy (abadi2016deep, ), and adversarial regularization (nasr2018machine, ). Salem et al. came up with the first effective defense mechanism against MIA (salem2018ml, ). Their approach included dropout and model stacking. In each training iteration of a fully connected neural network model, dropout is defined as the random deletion of a certain proportion of neuron connections. It can mitigate overfitting in deep neural networks, a contributing factor to MIA (salem2018ml, ). However, this technique works only when a neural network is targeted by the attack model. To work with other target models, they proposed another defense technique referred to as model stacking. The idea behind this defense is if distinct parts of the target model undergo training with different subsets of data, the overall model is expected to exhibit a lower tendency of overfitting. It can be achieved through the application of model stacking, one of the popular ensemble learning techniques. Differential privacy (DP) based techniques are also widely used to prevent privacy leakage by MIA (mattern2023membership, ), (li2021large, ). It includes data perturbation and output perturbation (hu2023defenses, ). Models facilitated with differential privacy employed with the stochastic gradient descent optimization algorithm (dwork2014algorithmic, ) can reduce empirical privacy leakages while ensuring comparable model utility in the non-DP environment (jagannatha2021membership, ). Another defense method against MIAs is to include regularization during the training of the model. Regularization refers to a set of techniques used to prevent overfitting and improve the generalization performance of an ML model. Label smoothing (szegedy2016rethinking, ) is one kind of regularization method that prevents overfitting of the ML model, which contributes to MIA (yeom2018privacy, ). Very few defense techniques have been proposed for LLMs (fu2023practical, ), (li2021large, ). Most of the existing defense techniques have been experimented on relatively small language models, such as test classifiers (wang2022analyzing, ), which are not evaluated for LLMs. Moreover, DP-based defenses may impair model utility. We argue that there is a pressing need for further research studies to develop effective defense techniques against MIA on LLMs.
@@
Defenses Against PII Leakage Attacks. To mitigate the leaking of personal information from PLMs due to memorization, there are several general techniques. During the preprocessing phase, the process of deduplication has the potential to significantly decrease the amount of memorized text in PLMs. Consequently, this results in a reduction of stored personal information within these models (lee2021deduplicating, ). Manually checking the vast training data for language models (LMs) is impractical. However, methods such as personal information or content identifying and filtering with restrictive terms of use can limit sensitive content (continella2017obfuscation, ), (ren2016recon, ). Deduplication at the document or paragraph level is common but may not eliminate repeated occurrences of sensitive information within a single document. Advanced strategies for deduplication and careful sourcing of training data are essential. Despite sanitization efforts, complete prevention of privacy leaks is challenging, making it a first line of defense rather than a foolproof measure. In training, following the process of (carlini2021extracting, ) and the implementation of (anil2021large, ), the differentially private stochastic gradient descent (DP-SGD) algorithm (abadi2016deep, ) can be employed to ensure privacy of training data during the training process (carlini2021extracting, ), (inan2021privacy, ). However, the DP-SGD-based method might not work efficiently as it has a significant computational cost and decreases the trained model utility (zanella2020analyzing, ). PII scrubbing filters dataset to eliminate PII from text (PIIsc, ), such as leveraging Named Entity Recognition (NER) (lample2016neural, ) to tag PII. Even though PII scrubbing methods can mitigate PII leakage risks, they face two critical challenges (lukas2023analyzing, ): (1) the effectiveness of PII scrubbing may be reduced to preserve the dataset utility, and (2) there is a risk of PII not being completely or accurately removed from the dataset. In downstream applications like dialogue systems (zhang2019dialogpt, ) and summarization models (hoang2019efficient, ), LMs undergo fine-tuning on task-specific data. While this process may lead to the LM “forgetting“ some memorized data from pretraining (mccloskey1989catastrophic, ), (ratcliff1990connectionist, ), it can still introduce privacy leaks if the task-specific data contains sensitive information. The aforementioned defense techniques mostly apply to the LMs. We have not yet identified specific techniques that are dedicated to the LLMs. According to some recent studies (li2023privacy, ) (yao2023survey, ), the existing strategies can be used in the LLM context as well. So far, there is a pressing need for more empirical evaluations to determine their effectiveness for LLMs. On top of that, there are no efficient defense techniques introduced to defend against a few attack methods, such as KART (nakamura2020kart, ), ProPILE (kim2023propile, ), and the recovery of forgotten PII by fine-tuning due to memorization (chen2023janus, ). Therefore, it requires in-depth studies and understanding to design effective defense techniques against PII attacks on LLMs.
@@ 7. Application-based Risks in LLMs
LLMs are emerging techniques with high potential for many applications. The security and privacy vulnerabilities of LLMs may raise serious concerns and risks in their real-world deployment with varying impacts on different application domains (tornede2023automl, ), (mozes2023use, ).
@@
Complicated Human-Interaction. LLM undergoes training on extensive text corpora and inherently possesses knowledge across diverse tasks. Carefully crafted prompts can potentially extract valuable and accurate knowledge from LLMs, which requires exploring and developing effective prompt engineering techniques (sorensen2022information, ). Despite the ideal scenario envisioning automated prompt generation through human-machine interaction, it is highly important to study the ethical issues and limitations in this approach (weidinger2021ethical, ). Consequently, a noteworthy concern emerges wherein the reliance on LLMs may potentially shift the entry barrier from coding and machine learning expertise to proficiency in prompt engineering.
@@
Misinformation and disinformation dissemination. LLMs are renowned for generating sound output that may incorporate hallucinated knowledge, posing challenges in distinguishing it from facts. This gives rise to concerns regarding potential adverse outcomes in LLM utilization, such as user misconfigurations leading to minimal run-time allocation or inappropriate decision-making in tasks like selecting search spaces for specific problems (ji2023survey, ). The deployment of LLMs also entails risks, including the creation of less informed users and the erosion of trust in shared information (weidinger2021ethical, ). Particularly in sensitive domains like legal or medical advice, misinformation can have serious consequences, potentially leading users to engage in illegal actions or follow detrimental instructions on medical conditions (pan2023risk, ), (weidinger2022taxonomy, ). Simultaneously, the intentional dissemination of fake news and disinformation carries severe implications, influencing public perception and decision-making processes, and contributing to societal discord (whitehouse2022evaluation, ). The dynamic nature of information dissemination in the digital age magnifies these risks, necessitating the development of robust fact-checking mechanisms, ethical guidelines for content generation, and responsible deployment practices for LLMs.
@@
Cybercrime and Social Issues. LLMs can potentially be used in various cybercrime (kshetri2023cybercrime, ), e.g., phishing (efficiently create targeted scam e-mails), malware, and hacking attacks (hackers have used ChatGPT to write malware codes). LLMs pose risks of perpetuating unfair discrimination and causing representational harm by reinforcing stereotypes and social biases. Harmful associations of specific traits with social identities may lead to exclusion or marginalization of individuals outside established norms (weidinger2021ethical, ). Additionally, toxic language generated by LLMs may incite hate or violence and cause serious offense. These risks are largely rooted in the selection of training corpora that include harmful language and disproportionately represent certain social identities.
@@
Transportation. In the transportation domain, studies reported that LLM can be biased (while doing accident report analysis), and inefficient for performing tasks in self-driving cars (tian2022federated, ). Furthermore, it might leak personal data from self-driving cars while doing accident report automation, and accident information extraction (trans, ). A framework has been proposed named VistaGPT to deal with the problems caused by information barriers from heterogeneity at both system and module levels in a wide range of heterogeneous vehicle automation systems (tian2023vistagpt, ). It leverages LLMs to create an automated composing platform to design end-to-end driving systems. This involves employing a “dividing and recombining” strategy to enhance the ability to generalize. To alleviate the issue of the long training time of LLMs with large datasets and high computing resource requirements, Meta-AI’s LLaMA focuses on fine-tuning offline pretrained LLMs to handle the transportation safety domain tasks. The main objective is to create a specialized LLM capable of generating an accurate, context-sensitive, and safety-aware model, that work effectively in traffic-related scenarios (zheng2023trafficsafetygpt, ).
@@
Healthcare and Medicine. The high risks associated with LLMs in the context of healthcare suggest that their integration into the healthcare system is presently inadvisable, as proposed by De et al. (de2023chatgpt, ). Models trained on extensive Internet data lacking rigorous filtering mechanisms may inadvertently incorporate misinformation, biased content, and harmful materials alongside accurate and fair information, thereby posing significant risks in healthcare applications. The potential consequences of erroneous treatment or medication recommendations by LLMs are particularly concerning. Moreover, the probabilistic nature of LLMs introduces variability in responses to the same task, giving rise to challenges in reliability and reproducibility that necessitate continuous human oversight. Privacy concerns, especially regarding sensitive health records, coupled with broader considerations such as AI ethics principles, safety, transparency, explainability, equity, and sustainability, further emphasize the need for caution in deploying LLMs within the healthcare domain, as discussed by Harrer et al. (harrer2023attention, ).
@@
Education. The use of LLMs, e.g., ChatGPT, in education is associated with significant drawbacks, particularly in fostering inaccurate concept learning and an inappropriate approach to education. ChatGPT, being a language model trained on diverse Internet data, may unintentionally propagate misinformation or present concepts with a lack of precision and educational rigor, for instance, scientific misconduct (mozes2023use, ). The excessive dependence on LLMs by both educators and learners can have serious adverse effects. Students engaging with ChatGPT may encounter misleading content or promote misconceptions, potentially compromising the quality of their learning experience. The absence of real-time fact-checking and the model’s susceptibility to biases and errors pose risks, potentially leading learners astray and impeding their overall educational progress. Consequently, caution is recommended when relying on ChatGPT as an educational tool without appropriate supervision and verification (senechal2023balancing, ), (milano2023large, ).
@@
Governance. The potential misuse of LLMs in governance for spear phishing presents significant cybersecurity challenges. Using GPT-4 as an illustrative example, personal information of British members of parliament (MP) was extracted from Wikipedia, and GPT-3.5 was utilized to generate biographies, which were then incorporated into phishing emails sent to official email addresses (hazell2023large, ). This highlights the risks associated with misinformation, biased content, and the utilization of LLMs in AI-based cyberattacks within governance. The leakage of confidential information through such attacks can pose severe consequences for national security. The generation of misinformation and hate speech by LLMs further emphasizes the existing challenges, underscoring the imperative need for robust safeguards and countermeasures to address the risks related to the usage of these models in governance settings (hazell2023large, ).
@@
Science. Hallucinations, biases, and paradigm shifts are pressing concerns of LLMs in the science domain. There is a risk of LLMs generating non-existent and false content. For instance, Meta developed an LLM named Galactica for reasoning scientific knowledge. That was reported to generate major flaws due to reproducing biases and presenting falsehoods (metaG, ). As a result, the model was shut down just after launching public access (heaven2022meta, ). Another concern lies in the involvement of LLMs in the scientific discovery process. It is challenging to interpret and understand LLMs due to their black-box nature, raising doubts about their reliability and trustworthiness in the science domain. For example, peer-review reports generated by LLMs may misinterpret research articles, which may impair the peer review quality (sci, ). Moreover, collaborating with LLMs won’t be fundamentally the same as collaborating with other researchers or experts in a corresponding field (binz2023should, ). Clear principles of using these LLMs and/or other AI tools in scientific explorations should be established to ensure transparency, fairness, and trustworthiness (binz2023should, ).

### v2

@@ 6.2. Defense Against Privacy Attacks on LLMs
Defense Against Gradient Leakage Attack. There are several mitigation strategies to defend against those gradient-based attack methods, e.g., TAG (Deng et al., 2021) and LAMP (Balunovic et al., 2022). They are random noise insertion to the gradients (Wei et al., 2020), differential privacy (DP) (Geyer et al., 2017; Abadi et al., 2016), and homomorphic encryption (Aono et al., 2017). DP is the most common and effective technique to mitigate the effect of gradient leakage attacks in DNNs. The fundamental idea of using DP is to add a controlled amount of noise to the model updates during training, which limits the model’s ability to memorize and reproduce original sequences from the training data (Inan et al., 2021b). Building upon prior research on vision model attacks (Zhu et al., 2019), (Wei et al., 2020), the defense mechanisms involving the addition of Gaussian or Laplacian noise to gradients and DP-SGD coupled with additional clipping (Abadi et al., 2016) can form an effective defense against gradient leakage attacks. For example, DP can efficiently mitigate the impact of TAB (Inan et al., 2021b) by reducing the number of unique training sequences leaked by Transformer-based language models. TAB depends on black-box access to the model, i.e., the model’s top-k predictions at each token position given an input prefix. Ensuring no access to the model’s underlying probability distributions through API hardening techniques (Inan et al., 2021b) may potentially mitigate the impact of such attacks. Previous studies (Geiping et al., 2020) have demonstrated that the success of gradient leakage attacks is highly dependent on the attacker’s ability to solve the gradient optimization problem over a loss function under the condition of non-zero gradients. Moreover, the attack strategy for successfully reconstructing private training data may differ between well-trained models and inadequately trained models. An effective mitigation technique could involve robust model optimization during training so that adversarial perturbations have minimum impact (Zhang et al., 2021a).
However, there is a significant shortcoming of using DP in model training. It may sacrifice the model’s utility to a certain extent. Trading-off preserving privacy and model utility may rise significant challenge for preventing gradient leakage attacks LLMs. Prior works explored various techniques, such as (Wu et al., 2023a), (Raeini, 2023), and (Huang et al., 2022) to defend against gradient leakage attacks in the language domain for small NLP models. Further research is required to develop defense mechanisms against gradient leakage attacks on LLMs.
@@
Defense Against Membership Inference Attack. In order to mitigate MIA in the language domain, several mechanisms are proposed, including dropout, auto de-identification (Vakili et al., 2022) model stacking, differential privacy (Abadi et al., 2016), and adversarial regularization (Nasr et al., 2018). Salem et al. came up with the first effective defense mechanism against MIA (Salem et al., 2018). Their approach included dropout and model stacking. In each training iteration of a fully connected neural network model, dropout is defined as the random deletion of a certain proportion of neuron connections. It can mitigate overfitting in DNNs, which is a contributing factor to MIA (Salem et al., 2018). However, this technique works only when a neural network is targeted by the attack model. To work with other target models, they proposed another defense technique referred to as model stacking. The idea behind this defense is if distinct parts of the target model undergo training with different subsets of data, the overall model is expected to exhibit a lower tendency of overfitting. It can be achieved through the application of model stacking, one of the popular ensemble learning techniques. To defend against black-box MIA on ML models, Jia et al. proposed MemGuard (Jia et al., 2019) which is essentially a noise perturbation mechanism for the predicted confidence score of the target model. It makes difficult for the attacker to infer whether a sample was part of the training data. However, this technique has been evaluated for image and numeric datasets. The performance of this technique on LMs/LLMs still remains unexplored, which necessitates further studies to evaluate whether it can effectively defend against MIA for LM/LLMs. Differential privacy (DP) based techniques are also widely used to prevent privacy leakage by MIA (Mattern et al., 2023), (Li et al., 2021c). It includes data perturbation and output perturbation (Hu et al., 2023b). Models facilitated with differential privacy employed with the stochastic gradient descent optimization algorithm (Dwork et al., 2014) can reduce empirical privacy leakages while ensuring comparable model utility in the non-DP environment (Jagannatha et al., 2021). Model pruning (Han et al., 2015), and knowledge distillation (Hinton, 2015) are employed to mitigate the impacts of preference-based MIA, such as PREMIA (Feng et al., 2024b). A recent framework, InferDPT (Tong et al., 2023), has been proposed to leverage black-box LLMs to facilitate privacy-preserving inference, which effectively integrates DP in text generation tasks. Another defense method against MIAs is to include regularization during the training of the model. Regularization refers to a set of techniques used to prevent overfitting and improve the generalization performance of an ML model. Label smoothing (Szegedy et al., 2016) is one kind of regularization method that prevents overfitting of the ML model, which contributes to MIA (Yeom et al., 2018). Very few defense techniques have been proposed for LLMs (Fu et al., 2023b), (Li et al., 2021c). Most of the existing defense techniques have been experimented on relatively small LMs, such as test classifiers (Wang et al., 2022b), which are not evaluated for LLMs. Moreover, DP-based defenses may impair model utility. We argue that there is a pressing need for further research studies to develop effective defense techniques against MIA on LLMs.
@@
Defenses Against PII Leakage Attacks. To mitigate the leaking of personal information from PLMs due to memorization (Carlini et al., 2021), there are several general techniques. During the pre-processing phase, the process of de-duplication and training data curation (Continella et al., 2017) has the potential to significantly decrease the amount of memorized text in PLMs. Consequently, this results in a reduction of stored personal information within these models (Lee et al., 2021). Prompt-tuning can also be a potential mitigation against memorization. The basic idea is to optimize prompts in the attack environment to evaluate the capability to extract the memorized content in a target model. DP can efficiently mitigate the impact of TAB (Inan et al., 2021b) and KART (Nakamura et al., 2020) by minimizing the number of unique training sequences leaked by Transformer-based language models. Both TAB and KART require the back-box access of the model, such as top-k predictions; therefore, restricting the model access by imposing API hardening technique (Inan et al., 2021b) can also be a potential mitigation technique for these kinds of attacks. DP can also mitigate the attacks that utilize hand-crafted prompts with true prefix (Nakka et al., 2024). PII masking (Mansfield et al., 2022) can also be a viable technique to defend against PII leakage attacks. Personal information identification and filtering methods, such as (Continella et al., 2017) and (Ren et al., 2016), may effectively reduce the number of training data samples extracted through PII leakage attacks, e.g., ProPILE (Kim et al., 2024). However, manually checking the vast training data for LLMs is challenging and labor-intensive. De-duplication at the document or paragraph level is common but may not eliminate repeated occurrences of sensitive information within a single document. Advanced strategies for de-duplication and careful sourcing of training data are essential. Despite sanitization efforts, complete prevention of privacy leaks is challenging, making it a first line of defense rather than a foolproof measure. In training, following the process of Carlini et al. (Carlini et al., 2021) and the implementation by Anil et al. (Anil et al., 2021), the deferentially private stochastic gradient descent (DP-SGD) algorithm (Abadi et al., 2016) can be employed to ensure privacy of training data during the training process (Carlini et al., 2021), (Inan et al., 2021a). However, the DP-SGD-based method might not work efficiently as it has a significant computational cost and decreases the trained model utility (Zanella-Béguelin et al., 2020). PII scrubbing filters dataset to eliminate PII from text (develop.sentry.dev, 2023), such as leveraging Named Entity Recognition (NER) (Lample et al., 2016) to tag PII. Even though PII scrubbing methods can mitigate PII leakage risks, they face two critical challenges (Lukas et al., 2023): (1) the effectiveness of PII scrubbing may be reduced to preserve the dataset utility, and (2) there is a risk of PII not being completely or accurately removed from the dataset. In downstream applications like dialogue systems (Zhang et al., 2019) and summarization models (Hoang et al., 2019), LMs undergo fine-tuning on task-specific data. While this process may lead to the LM “forgetting“ some memorized data from pre-training (McCloskey and Cohen, 1989), (Ratcliff, 1990), it can still introduce privacy leaks if the task-specific data contains sensitive information.
The aforementioned defense techniques mostly apply to the LMs. To date, we have not yet identified specific techniques that are dedicated to the LLMs. While according to some recent studies ((Li et al., 2023a), (Yao et al., 2024)), the existing strategies can be used in the LLM context as well. So far, there is a pressing need for more empirical evaluations to determine their effectiveness for LLMs. On top of that, there are no efficient defense techniques introduced to defend against a few attack methods, such as KART (Nakamura et al., 2020), ProPILE (Kim et al., 2024), and the recovery of forgotten PII by fine-tuning due to memorization (Chen et al., 2023). Therefore, it requires in-depth studies and understanding to design effective defense techniques against PII attacks on LLMs.
@@ 7. Application-based Risks in LLMs
LLMs are emerging techniques with high potential for many applications. The security and privacy vulnerabilities of LLMs may raise serious concerns and risks in their real-world deployment with varying impacts on different application domains (Tornede et al., 2023), (Mozes et al., 2023).
@@
Complicated Human-Interaction. LLM undergoes training on extensive text corpora and inherently possesses knowledge across diverse tasks. Carefully crafted prompts can potentially extract valuable and accurate knowledge from LLMs, which requires exploring and developing effective prompt engineering techniques (Sorensen et al., 2022). Despite the ideal scenario of envisioning automated prompt generation through human-machine interaction, it is highly important to study the ethical issues and limitations of this approach (Weidinger et al., 2021). Consequently, a noteworthy concern emerges wherein the reliance on LLMs may potentially shift the entry barrier from coding and machine learning expertise to proficiency in prompt engineering.
@@
Hallucination, misinformation and disinformation dissemination. LLMs are renowned for generating sound output that may incorporate hallucinated knowledge, falsification (Jansson et al., 2021), misinterpretation (Amaratunga, 2023), biasness (Dai et al., 2024) which pose challenges in distinguishing it from facts. This gives rise to concerns regarding potential adverse outcomes in LLM utilization, such as user misconfigurations leading to minimal run-time allocation or inappropriate decision-making in tasks like selecting search spaces for specific problems (Ji et al., 2023a). Catastrophic forgetting is a problem where a neural network forgets information it previously learned after being trained on a new task. In LLM fine-tuning, it might deteriorate the performance at significant extent (Luo et al., 2023b). The deployment of LLMs also entails risks, including the creation of less informed users and the erosion of trust in shared information (Weidinger et al., 2021). Particularly in sensitive domains like legal or medical advice, misinformation can have serious consequences, potentially leading users to engage in illegal actions or follow detrimental instructions on medical conditions (Pan et al., 2023), (Weidinger et al., 2022). Simultaneously, the intentional dissemination of fake news and disinformation carries severe implications, influencing public perception and decision-making processes, and contributing to societal discord (Whitehouse et al., 2022). LLM-generated misinformation can cause harm to many important sectors of society such as politics (Pierri et al., 2023), finance (Rangapur et al., 2023), healthcare (Perlis et al., 2023), and so on (Chen and Shu, 2023b). The dynamic nature of information dissemination in the digital age magnifies these risks, necessitating the development of robust fact-checking mechanisms, ethical guidelines for content generation, and responsible deployment practices for LLMs. Hallucination mitigation can be applied at both training (e.g., data curation (Yan et al., 2024a) and knowledge enhancement (Hu et al., 2023a)) and inference stages (e.g., uncertainty measurement (Huang et al., 2023a), knowledge retrieval (Feng et al., 2024a), and self-familiarity (Luo et al., 2023a)). Recent studies have introduced several countermeasures, such as the chain of verification (COVE (Dhuliawala et al., 2023)) and self-reflection (Ji et al., 2023b), to detect LLM-generated hallucinations and misinformation (Chen and Shu, 2023a; Chern et al., 2023). These techniques offer potential remedies for the issues mentioned above; however, further exploration is necessary to fully address these challenges. To address catastrophic forgetting, learning rate scheduling (LR-Adjust (Winata et al., 2023; Jin et al., 2023b; Wu and Liu, 2023)) serves as a viable mitigation technique. Self-Synthesized Rehearsal (Huang et al., 2024) (SSR) is another approach to mitigate this issue. SSR employs a base LLM to generate synthetic instances for in-context learning, which is then refined by the latest LLM, preserving its learning ability. High-quality synthetic outputs are chosen for future rehearsals to mitigate catastrophic forgetting.
@@
Cybercrime and Social Issues. LLMs can potentially be used in various cybercrime (Kshetri, 2023), e.g., phishing (efficiently create targeted scam e-mails), malware, and hacking attacks (hackers have used ChatGPT to write malware codes). LLMs pose risks of perpetuating unfair discrimination and causing representational harm by reinforcing stereotypes and social biases. Harmful associations of specific traits with social identities may lead to exclusion or marginalization of individuals outside established norms (Weidinger et al., 2021). Additionally, toxic language generated by LLMs may incite hate or violence and cause serious offense. These risks are largely rooted in the selection of training corpora that include harmful language and disproportionately represent certain social identities.
@@
Transportation. In the transportation domain, studies reported that LLM can be biased (while doing accident report analysis), and inefficient for performing tasks in self-driving cars (Tian et al., 2022). Furthermore, it might leak personal data from self-driving cars while doing accident report automation, and accident information extraction (Zheng et al., 2023c). A framework has been proposed named VistaGPT to deal with the problems caused by information barriers from heterogeneity at both system and module levels in a wide range of heterogeneous vehicle automation systems (Tian et al., 2023). It leverages LLMs to create an automated composing platform to design end-to-end driving systems. This involves employing a “dividing and recombining” strategy to enhance the ability to generalize. To alleviate the issue of the long training time of LLMs with large datasets and high computing resource requirements, Meta-AI’s Llama focuses on fine-tuning offline pre-trained LLMs to handle the transportation safety domain tasks. The main objective is to create a specialized LLM capable of generating an accurate, context-sensitive, and safety-aware model, that work effectively in traffic-related scenarios (Zheng et al., 2023a).
@@
Healthcare and Medicine. The high risks associated with LLMs in the context of healthcare suggest that their integration into the healthcare system is presently inadvisable, as proposed by De et al. (De Angelis et al., 2023). Models trained on extensive Internet data lacking rigorous filtering mechanisms may inadvertently incorporate misinformation, biased content, and harmful materials alongside accurate and fair information, thereby posing significant risks in healthcare applications. The potential consequences of erroneous treatment or medication recommendations by LLMs are particularly concerning. Moreover, the probabilistic nature of LLMs introduces variability in responses to the same task, giving rise to challenges in reliability and reproducibility that necessitate continuous human oversight. Privacy concerns, especially regarding sensitive health records, coupled with broader considerations such as AI ethics principles, safety, transparency, explainability, equity, and sustainability, further emphasize the need for caution in deploying LLMs within the healthcare domain, as discussed by Harrer et al. (Harrer, 2023).
@@
Education. The use of LLMs, e.g., ChatGPT, in education is associated with significant drawbacks, particularly in fostering inaccurate concept learning and an inappropriate approach to education. ChatGPT, being a language model trained on diverse Internet data, may unintentionally propagate misinformation or present concepts with a lack of precision and educational rigor, for instance, scientific misconduct (Mozes et al., 2023). The excessive dependence on LLMs by both educators and learners can have serious adverse effects. Students engaging with ChatGPT may encounter misleading content or promote misconceptions, potentially compromising the quality of their learning experience. The absence of real-time fact-checking and the model’s susceptibility to biases and errors pose risks, potentially leading learners astray and impeding their overall educational progress. Consequently, caution is recommended when relying on ChatGPT as an educational tool without appropriate supervision and verification (Senechal et al., 2023), (Milano et al., 2023).
@@
Governance. The potential misuse of LLMs in governance for spear phishing presents significant cybersecurity challenges. Using GPT-4 as an illustrative example, personal information of British members of parliament (MP) was extracted from Wikipedia, and GPT-3.5 was utilized to generate biographies, which were then incorporated into phishing emails sent to official email addresses (Hazell, 2023). This highlights the risks associated with misinformation, biased content, and the utilization of LLMs in AI-based cyberattacks within governance. The leakage of confidential information through such attacks can pose severe consequences for national security. The generation of misinformation and hate speech by LLMs further emphasizes the existing challenges, underscoring the imperative need for robust safeguards and countermeasures to address the risks related to the usage of these models in governance settings (Hazell, 2023).
@@
Science. Hallucinations, biases, and paradigm shifts are pressing concerns of LLMs in the science domain. There is a risk of LLMs generating non-existent and false content. For instance, Meta developed an LLM named Galactica for reasoning scientific knowledge. That was reported to generate major flaws due to reproducing biases and presenting falsehoods (Conversation, 2022). As a result, the model was shut down just after launching public access (Heaven, 2022). Another concern lies in the involvement of LLMs in the scientific discovery process. It is challenging to interpret and understand LLMs due to their black-box nature, raising doubts about their reliability and trustworthiness in the science domain. For example, peer-review reports generated by LLMs may misinterpret research articles, which may impair the peer review quality (Abeba Birhane, 2023). Moreover, collaborating with LLMs won’t be fundamentally the same as collaborating with other researchers or experts in a corresponding field (Binz et al., 2023). Clear principles of using these LLMs and/or other AI tools in scientific explorations should be established to ensure transparency, fairness, and trustworthiness (Binz et al., 2023).